# A Closer Look at AUROC and AUPRC under Class Imbalance

### v3

@@ 7 Conclusion
This study rigorously interrogates the pervasive assumption within the machine learning community that AUPRC is a more appropriate evaluation metric than AUROC in class-imbalanced settings. Our empirical analyses, along with an exhaustive literature review, have revealed several important findings that critically challenge this belief. In particular, we show that while optimizing for AUROC equates to minimizing the model’s FPR in an unbiased manner over positive sample scores, optimizing for AUPRC equates to minimizing the FPR specifically for regions where the model outputs higher scores relative to lower scores. We further show both theoretically and empirically over synthetic and real-world fairness datasets that AUPRC can be an explicitly discriminatory metric in that it favors higher-prevalence subgroups.
@@
Broader Impact and Ethical Considerations
@@
While we use the COMPAS dataset for recividism prediction in this work, we recognize the many societal issues with automated predictions of recidivism (Dressel & Farid, 2018). We utilize this dataset as it is a commonly used dataset in the fairness literature, but do not advocate for deployment of these models in any way.
@@
Acknowledgements
MBAM is supported by a Berkowitz Postdoctoral Fellowship. JG is funded by the National Institute of Health through DS-I Africa U54 TW012043-01 and Bridge2AI OT2OD032701.
Appendix A Code Availability
Appendix B Notation
@@
We define random variable $\mathsf{p}=f_{\boldsymbol{\theta}}(\mathsf{x})$ to be the distribution of scores output by the model over input samples. Throughout the paper, $\boldsymbol{\theta}$ may be omitted if it is clear from context. We will occasionally also use the notation $\mathsf{p}_{+}$ and $\mathsf{p}_{-}$ to reflect the conditional distributions of model outputs conditioned on the label being 1 or 0, respectively:
$\mathsf{p}_{+}$ $=f(\mathsf{x})|\mathsf{y}=1$
$\mathsf{p}_{-}$ $=f(\mathsf{x})|\mathsf{y}=0.$
@@
$=P_{\mathsf{p}|\mathsf{y}=1}(p>\tau)$
$=P(p_{+}>\tau)$
@@
$=P_{\mathsf{p}|\mathsf{y}=0}(p>\tau)$
$=P(p_{-}>\tau)$
@@
$=P_{\mathsf{y}|\mathsf{p}>\tau}(y=1)$
@@
Appendix C Proof of Theorem 1
Recall that all notation is defined formally in Appendix B.
@@ Appendix C Proof of Theorem 1
$=\mathbb{E}_{\mathsf{p}_{+}}\left[g\right]$
So, $\mathrm{AUROC}=1-\mathbb{E}_{\mathsf{p}_{+}}\left[\mathrm{FPR}\right]$ $\mathrm{AUPRC}=\mathbb{E}_{\mathsf{p}_{+}}\left[\mathrm{Prec}\right]$. To further simplify, we expand $\mathrm{Prec}$ via Bayes rule:
$\mathrm{Prec}$ $=1-P_{\mathsf{y}|\mathsf{p}>\tau}(y=0)$
$=1-\underbrace{P_{\mathsf{p}|\mathsf{y}=0}(p>\tau)}_{\mathrm{FPR}(\tau)}\frac{P_{\mathsf{y}}(y=0)}{P_{\mathsf{p}}(p>\tau)}$
@@
$\mathrm{AUROC}$ $=1-\mathbb{E}_{p_{+}}\left[\mathrm{FPR}\right]$
$\mathrm{AUPRC}$ $=\mathbb{E}_{p_{+}}\left[\mathrm{Prec}\right]$
$=1-P_{\mathsf{y}}(y=0)\mathbb{E}_{p_{+}}\left[\frac{\mathrm{FPR}}{P_{\mathsf{p}}(p>p_{+})}\right],$
@@
Synthetic validation of Theorem 1 can also be found in our public code.
Appendix D Proof of Theorem 2
@@
Appendix E Proof of Theorem 3
@@
Let a model $f$ be perfectly calibrated and yield score distributions for positive and negative samples from $p_{+}$ and $p_{-}$. Then $p_{+}(t)=\frac{t}{1-t}\frac{p_{\mathsf{y}}(0)}{p_{\mathsf{y}}(1)}p_{-}(t)$
@@
$p_{+}(t)$ $=P_{\mathsf{p}|\mathsf{y}=1}(t)$
$=\frac{P_{\mathsf{y}|\mathsf{p}=t}(1)P_{\mathsf{p}}(t)}{p_{\mathsf{y}}(1)}$
$=t\frac{p_{\mathsf{y}}(1)p_{+}(t)+p_{\mathsf{y}}(0)p_{-}(t)}{p_{\mathsf{y}}(1)}$
$=tp_{+}(t)+t\frac{p_{\mathsf{y}}(0)}{p_{\mathsf{y}}(1)}p_{-}(t).$
Thus, $p_{+}(t)=\frac{t}{1-t}\frac{p_{\mathsf{y}}(0)}{p_{\mathsf{y}}(1)}p_{-}(t)$ as desired. ∎
@@
Let a model $f$ be perfectly calibrated and yield score distributions for positive and negative samples from $p_{+}$ and $p_{-}$, with overall distribution given by $p(t)=p_{\mathsf{y}}(1)p_{+}(t)+p_{\mathsf{y}}(0)p_{-}(t)$. Then for all $\tau\in(0,1)$, $\mathrm{FR}(f,\tau)\leq\frac{p_{\mathsf{y}}(1)}{\tau}$.
@@
$\mathrm{FR}(f,\tau)$ $=\int_{\tau}^{1}p_{\mathsf{y}}(1)p_{+}(t)+p_{\mathsf{y}}(0)p_{-}(t)dt$
$=\int_{\tau}^{1}p_{\mathsf{y}}(1)p_{+}(t)+p_{\mathsf{y}}(1)\frac{1-t}{t}p_{+}(t)dt$
@@
$\leq p_{\mathsf{y}}(1)\cdot\frac{1}{\tau}\cdot 1$
$=\frac{p_{\mathsf{y}}(1)}{\tau}.$
@@
Appendix F Details for Synthetic Experiments
F.1 Sampling a random model with a given AUROC
@@
    If desired, the output scores can further be scaled to have expectation given by the dataset’s prevalence or can be adjusted via a calibration method to be calibrated given the assigned labels. Both procedures can be done without affecting the AUROC. Note that as any calibrated model will have expected probability given by the label’s prevalence (See Appendix F.2), the former condition is strictly weaker than the latter.
F.2 Calibration includes prevalence matching
Let $\mathsf{p}$ be a random variable describing the probabilities output by the model over the input distribution defined by the data generative function. If a model is calibrated, this means that $P_{\mathsf{y}|\mathsf{p}}(y=1|p=q)=q$ — that the probability that the label for a given point is 1 is given precisely by the models output probability for that sample. With that in mind, we have:
@@ F.2 Calibration includes prevalence matching
F.3 Details on optimization procedures
@@
Appendix G Additional Details on Real World Experiments
G.1 Dataset Details
@@
  • adult (Asuncion & Newman, 2007): The UCI Adult dataset, where the goal is to predict whether an individual’s income is >\$$>\$50k$.
  • compas (Angwin et al., 2022): The task to predict two-year recidivism. We only select samples belonging to “African-American” and “Caucasian”, leading to a binary race variable.
  • lsac (Wightman, 1998): The task is to predict whether a law school applicant will pass the bar. We only select samples belonging to White and Black applicants.
  • mimic (Johnson et al., 2016): We use the in-hospital mortality task proposed by Harutyunyan et al. (2019), where the goal is to predict whether a patient will die in the ICU given labs and vitals from the first 48 hours of their hospital stay. We only select samples belonging to White and Black patients.
@@
Table 1: Dataset statistics for the four binary classification datasets used in this study. Note that $n$ refers to the number of samples after balancing by the corresponding attribute.
@@ G.2 Hyperparameter Grid
G.3 Additional Results
Figure 7: Spearman’s $\rho$ between the test-set signed AUROC gap versus the validation set overall AUPRC, and the AUROC gap versus the overall AUROC. Numbers in parentheses are the prevalence ratios between the two groups for the particular attribute, and datasets are sorted by this quantity. Error bars are 95% confidence intervals from five different random data splits.
Figure 8: Correlation between the prevalence ratio, and the difference between the Spearman’s $\rho$ of the AUROC gap versus AUROC and the AUROC gap versus AUPRC. Each point represents a dataset and attribute combination. This correlation itself has a Spearman’s $\rho$ of $0.714$ (p = 0.047).
Appendix H Literature Review Methodology
H.1 Paper Acquisition
The initial phase of our comprehensive literature search involved the acquisition of datasets from both the Arxiv repository and NeurIPS conference proceedings. Utilizing the RedPajama dataset available through Hugging Face for Arxiv papers, and a custom scraper for NeurIPS papers, we specifically targeted these sources for our literature review. The Arxiv dataset, approximately 93.8 GB in size, encompassed over 1.5 million texts in JSONL format. For NeurIPS, we developed a script to scrape conference papers from 1987 to 2019 (9680 texts), aiming to enrich our search.
H.2 Keyword-Driven Filtering Process
@@
H.3 AI-Assisted Screening and Refinement
@@ H.4 Manual Review
H.4.1 Final Papers
After manual review, we identified 128 papers that make or reference some version of the claim that “AUPRC is better than AUROC in cases of class imbalance.” (Cook & Ramadas, 2020; Leisman, 2018; Yang et al., 2015; Gaudreault et al., 2021; Albora & Zaccaria, 2022; Lim & van der Schaar, 2018; Liu et al., 2023; Randl et al., 2023; Tusfiqur et al., 2022; Piermarini et al., 2023; Zhang & Bondell, 2018; Weiss & Tonella, 2021; Afanasiev et al., 2021; Li et al., 2022; Torfi et al., 2022; Wu et al., 2020; Miao & Zhu, 2022; Navarro et al., 2022; Cho et al., 2021; Wagner et al., 2023; Isupova et al., 2017; Sarvari et al., 2021; Hiri et al., 2022; Herbach, 2021; Si & Roberts, 2021; Narayanan et al., 2022; Li et al., 2020; Lee et al., 2013; Rayhan et al., 2017; Kyono et al., 2018; Adler, 2021; Seo et al., 2021; Hong et al., 2019; Hagedoorn & Spanakis, 2017; Yang et al., 2022a; Babaei et al., 2021; Garcin & Stéphan, 2021; Mehboudi et al., 2022; Yang et al., 2022b; Shen & Kursun, 2021; Muthukrishna et al., 2019; Deng et al., 2023; Yang, 2021; Harer et al., 2018; Meister et al., 2022; Skarding et al., 2021; Alvarez et al., 2022; Zou et al., 2022; Mangolin et al., 2022; Mosteiro et al., 2021; Hashemi et al., 2018; Lee et al., 2020; Zavrtanik et al., 2021; Showalter & Wu, 2019; Cranmer & Desmarais, 2016; Bryan & Moriano, 2023; Zhang et al., 2017; Domingues et al., 2020; Markdahl et al., 2017; Fu et al., 2021; Pang et al., 2023; Rezvani et al., 2021; Ozyegen et al., 2022; Prapas et al., 2023; Rayhan et al., 2020; Thambawita et al., 2020; Shukla & Marlin, 2019; Blevins et al., 2021; Vijayan et al., 2017; Budka et al., 2021; Hsu et al., 2020; Smith et al., 2023; Choi et al., 2018; Ju et al., 2018; Pashchenko et al., 2018; Chu et al., 2018; Silva et al., 2022; Bach Nguyen et al., 2022; Deshwar et al., 2015; Brophy & Lowd, 2020; Mayaki & Riveill, 2022; Mongia et al., 2021; Tiulpin et al., 2019; Romero et al., 2022; Rubin et al., 2012; Schwarz et al., 2021; Lyu et al., 2021; Lopez-Martinez et al., 2022; Ahmed & Courville, 2020; Gong et al., 2021; Zhang et al., 2021; Shukla & Marlin, 2018; Lund et al., 2019; Ma et al., 2022; Ruff et al., 2021; Lei Ba et al., 2015; Chakraborty et al., 2023; Rajabi & He, 2021; Newby et al., 2022; Axelrod & Gomez-Bombarelli, 2023; Kim et al., 2022; Ando & Huang, 2017; Stolman et al., 2022; Mosquera et al., 2022; Kulkarni et al., 2021; Won et al., 2019; Stephenson et al., 2022; Srivastava et al., 2019; Moor et al., 2019; Danesh Pazho et al., 2023; Kiran et al., 2018; Steinbuss & Böhm, 2021; Ma et al., 2020; Karadzhov et al., 2022; Ding et al., 2018; Mousavian et al., 2016; Rayhan et al., 2017; Vens et al., 2008; Rohani & Eslahchi, 2019; López et al., 2013; Sahiner et al., 2017; Rao et al., 2022; Hibshman & Weninger, 2023; Ntroumpogiannis et al., 2023; Weiss & Tonella, 2023; Hall et al., 2023; Goyal & Khiari, 2020; Boyd et al., 2013).
All papers identified, manual screening results, and extracted quotes can be found here: https://docs.google.com/spreadsheets/d/1NjDpwoj_8EkIwtGZzwM6w2nbst-LlGJPAqUcVINmPEk/edit?usp=sharing.

### v4

@@ 7 Conclusion
This study interrogates the pervasive assumption within the ML community that AUPRC is a better evaluation metric than AUROC in class-imbalanced settings. Our empirical analyses and literature review reveal several concrete findings that challenge this notion. In particular, we show that while optimizing for AUROC equates to minimizing the model’s FPR in an unbiased manner over positive sample scores, optimizing for AUPRC equates to minimizing the FPR specifically for regions where the model outputs higher scores relative to lower scores. Further, we show both theoretically and empirically over synthetic and real-world fairness datasets that AUPRC can be an explicitly discriminatory metric through favoring higher-prevalence subgroups.
@@
Acknowledgements
MBAM gratefully acknowledges support from a Berkowitz Postdoctoral Fellowship. JG is funded by the National Institute of Health through DS-I Africa U54 TW012043-01 and Bridge2AI OT2OD032701.
Appendix A Broader Impact and Ethical Considerations
@@
While we use the COMPAS dataset for recividism prediction in this work, we recognize the many societal issues with automated predictions of recidivism [93]. We utilize this dataset as it is a commonly used dataset in the fairness literature, but do not advocate for deployment of these models in any way.
@@
Appendix B Code Availability
Appendix C Notation
@@
We define random variable $\mathsf{s}=f_{\boldsymbol{\theta}}(\mathsf{x})$ to be the distribution of scores output by the model over input samples. Throughout the paper, $\boldsymbol{\theta}$ may be omitted if it is clear from context. We will occasionally also use the notation $\mathsf{s}_{+}$ and $\mathsf{s}_{-}$ to reflect the conditional distributions of model outputs conditioned on the label being 1 or 0, respectively:
$\mathsf{s}_{+}$ $=f(\mathsf{x})|\mathsf{y}=1$
$\mathsf{s}_{-}$ $=f(\mathsf{x})|\mathsf{y}=0.$
@@
$=P_{\mathsf{x}|\mathsf{y}=1}(f(x)>\tau)$
$=P_{\mathsf{s}|\mathsf{y}=1}(s>\tau)$
$=P(s_{+}>\tau)$
@@
$=P_{\mathsf{x}|\mathsf{y}=0}(f(x)>\tau)$
$=P_{\mathsf{s}|\mathsf{y}=0}(s>\tau)$
$=P(s_{-}>\tau)$
@@
$=P_{\mathsf{y}|f(\mathsf{x})>\tau}(y=1)$
$=P_{\mathsf{y}|\mathsf{s}>\tau}(y=1)$
@@
Appendix D Proof of Theorem 1
Recall that all notation is defined formally in Appendix C.
@@ Appendix D Proof of Theorem 1
$=\mathbb{E}_{\mathsf{s}_{+}}\left[g\right]$
So, $\mathrm{AUROC}=1-\mathbb{E}_{\mathsf{s}_{+}}\left[\mathrm{FPR}\right]$ $\mathrm{AUPRC}=\mathbb{E}_{\mathsf{s}_{+}}\left[\mathrm{Prec}\right]$. To further simplify, we expand $\mathrm{Prec}$ via Bayes rule:
$\mathrm{Prec}$ $=1-P_{\mathsf{y}|\mathsf{s}>\tau}(y=0)$
$=1-\underbrace{P_{\mathsf{s}|\mathsf{y}=0}(s>\tau)}_{\mathrm{FPR}(\tau)}\frac{P_{\mathsf{y}}(y=0)}{P_{\mathsf{s}}(s>\tau)}$
@@
$\mathrm{AUROC}(f)$ $=1-\mathbb{E}_{t\sim\mathsf{s}_{+}}\left[\mathrm{FPR}(f,t)\right]$
$=1-\mathbb{E}_{t\sim f(\mathsf{x})|\mathsf{y}=1}\left[\mathrm{FPR}(f,t)\right]$
$\mathrm{AUPRC}(f)$ $=\mathbb{E}_{t\sim\mathsf{s}_{+}}\left[\mathrm{Prec}(f,t)\right]$
$=1-P_{\mathsf{y}}(y=0)\mathbb{E}_{t\sim s_{+}}\left[\frac{\mathrm{FPR}(f,t)}{P_{s\sim\mathsf{s}}(s>t)}\right]$
$=1-P_{\mathsf{y}}(y=0)\mathbb{E}_{t\sim f(\mathsf{x})|\mathsf{y}=1}\left[\frac{\mathrm{FPR}(f,t)}{P_{\mathsf{x}}(f(x)>t)}\right]$
@@
Synthetic validation of Theorem 1 can also be found in our public code. Note that this formulation of AUPRC reflects earlier, different formulations of AUPRC, such as those found in the AUPRC optimization literature [409].
Appendix E Proof of Theorem 2
@@
Appendix F Proof of Theorem 3
@@
Let a model $f$ be perfectly calibrated and yield score distributions for positive and negative samples from probability density functions $p_{+}$ and $p_{-}$. Then $p_{+}(t)=\frac{t}{1-t}\frac{P_{\mathsf{y}}(y=0)}{P_{\mathsf{y}}(y=1)}p_{-}(t)$
@@
$p_{+}(t)$ $=P_{\mathsf{s}|\mathsf{y}=1}(s=t)$
$=\frac{P_{\mathsf{y}|\mathsf{s}=t}(y=1)p_{\mathsf{s}}(t)}{P_{\mathsf{y}}(y=1)}$
$=t\frac{P_{\mathsf{y}}(y=1)p_{+}(t)+P_{\mathsf{y}}(y=0)p_{-}(t)}{P_{\mathsf{y}}(y=1)}$
$=tp_{+}(t)+t\frac{P_{\mathsf{y}}(y=0)}{P_{\mathsf{y}}(y=1)}p_{-}(t).$
Thus, $p_{+}(t)=\frac{t}{1-t}\frac{P_{\mathsf{y}}(y=0)}{P_{\mathsf{y}}(y=1)}p_{-}(t)$ as desired. ∎
@@
Let a model $f$ be perfectly calibrated and yield score distributions for positive and negative samples from probability density functions $p_{+}$ and $p_{-}$, with overall distribution given by $p(t)=P_{\mathsf{y}}(y=1)p_{+}(t)+P_{\mathsf{y}}(y=0)p_{-}(t)$. Then for all $\tau\in(0,1)$, $\mathrm{FR}(f,\tau)\leq\frac{P_{\mathsf{y}}(y=1)}{\tau}$.
@@
$\mathrm{FR}(f,\tau)$ $=\int_{\tau}^{1}P_{\mathsf{y}}(y=1)p_{+}(t)+P_{\mathsf{y}}(y=0)p_{-}(t)dt$
$=\int_{\tau}^{1}P_{\mathsf{y}}(y=1)p_{+}(t)+P_{\mathsf{y}}(y=1)\frac{1-t}{t}p_{+}(t)dt$
@@
$\leq P_{\mathsf{y}}(y=1)\cdot\frac{1}{\tau}\cdot 1$
$=\frac{P_{\mathsf{y}}(y=1)}{\tau}.$
@@
Appendix G Details for Synthetic Experiments
G.1 Sampling a random model with a given AUROC
@@
    If desired, the output scores can further be scaled to have expectation given by the dataset’s prevalence or can be adjusted via a calibration method to be calibrated given the assigned labels. Both procedures can be done without affecting the AUROC. Note that as any calibrated model will have expected probability given by the label’s prevalence (See Appendix G.2), the former condition is strictly weaker than the latter.
The procedure outlined above guarantees that, in expectation, the AUROC of the generated set of scores and labels will be precisely the target AUROC. However, if you apply this procedure indpendently across different sample subpopulations, this guarantee can only be applied on each subpopulation individually, and not necessarily on the overall population due to the unspecified xAUC term. However, in practice, for the experiments we ran here, that impact neither meaningfully impacts our experiments nor were the joint AUROCs sufficiently different from the target AUROC to warrant a more complex methodology.
G.2 Calibration includes prevalence matching
Let $\mathsf{p}$ be a random variable describing the probabilities (not scores) output by the model over the input distribution defined by the data generative function. If a model is calibrated, this means that $P_{\mathsf{y}|\mathsf{p}}(y=1|p=q)=q$ — that the probability that the label for a given point is 1 is given precisely by the models output probability for that sample. With that in mind, we have:
@@ G.2 Calibration includes prevalence matching
G.3 Details on optimization procedures
@@
Appendix H Additional Details on Real World Experiments
H.1 Dataset Details
@@
  • adult [17]: The UCI Adult dataset, where the goal is to predict whether an individual’s income is >\$$>\$50k$.
  • compas [14]: The task to predict two-year recidivism. We only select samples belonging to “African-American” and “Caucasian”, leading to a binary race variable.
  • lsac [413]: The task is to predict whether a law school applicant will pass the bar. We only select samples belonging to White and Black applicants.
  • mimic [178]: We use the in-hospital mortality task proposed by [138], where the goal is to predict whether a patient will die in the ICU given labs and vitals from the first 48 hours of their hospital stay. We only select samples belonging to White and Black patients.
@@
Table 1: Dataset statistics for the four binary classification datasets used in this study. Note that $n$ refers to the number of samples after balancing by the corresponding attribute. Here, “Prevalence (Higher)” refers to the rate at which the prediction label $y=1$ for the subpopulation with a higher such rate, and “Prevalence (Lower)” refers to the same rate but over the subpopulation of the dataset with a lower rate of $y=1$.
@@ H.2 Hyperparameter Grid
H.3 Additional Results
All raw AUC results can be found in the code repository for these experiments, at https://github.com/hzhang0/auc_bias. Additionally, summarized results in different views can be found in Figures 7 and 8.
Figure 7: Spearman’s $\rho$ between the test-set signed AUROC gap versus the validation set overall AUPRC, and the AUROC gap versus the overall AUROC. Numbers in parentheses are the prevalence ratios between the two groups for the particular attribute, and datasets are sorted by this quantity. Error bars are 95% confidence intervals from 20 different random data splits.
Figure 8: Correlation between the prevalence ratio, and the difference between the Spearman’s $\rho$ of the AUROC gap versus AUROC and the AUROC gap versus AUPRC. Each point represents a dataset and attribute combination. This correlation itself has a Spearman’s $\rho$ of $0.905$ (p = 0.002).
Appendix I Literature Review Methodology
I.1 Paper Acquisition
The initial phase of our comprehensive literature search involved the acquisition of datasets from both the ArXiv preprint server (through the RedPajama dataset on Hugging Face), as well as from a subset of years of NeurIPS, ICML, ICLR, ACL, and CVPR conference proceedings (all scraped manually). The ArXiv dataset, approximately 93.8 GB in size, encompassed over 1.5 million texts in JSONL format. For NeurIPS, we developed a script to scrape conference papers from 1987 to 2019 (9680 texts), aiming to enrich our search. Other venues contributed fewer papers to our assessment process.
I.2 Keyword-Driven Filtering Process
@@
I.3 AI-Assisted Screening and Refinement
@@ I.4 Manual Review
I.4.1 Final Papers
After manual review, we identified 424 papers that make or reference some version of the claim that “AUPRC is better than AUROC in cases of class imbalance.” [78, 212, 428, 116, 8, 223, 229, 318, 393, 308, 450, 407, 4, 217, 391, 417, 261, 282, 70, 399, 169, 343, 153, 148, 361, 281, 219, 206, 204, 2, 350, 156, 134, 425, 22, 115, 257, 429, 353, 279, 85, 424, 137, 258, 367, 10, 454, 251, 275, 140, 208, 438, 356, 79, 52, 444, 90, 252, 111, 296, 325, 293, 312, 321, 387, 358, 41, 397, 54, 159, 369, 71, 180, 301, 74, 363, 24, 86, 49, 253, 270, 390, 329, 333, 348, 238, 233, 6, 124, 448, 359, 236, 243, 335, 210, 58, 317, 284, 21, 192, 12, 379, 274, 202, 415, 377, 373, 271, 82, 193, 376, 242, 183, 88, 276, 322, 395, 327, 239, 338, 319, 149, 288, 408, 136, 128, 46, 163, 132, 28, 207, 411, 302, 412, 227, 101, 221, 145, 328, 277, 234, 73, 60, 418, 92, 20, 198, 103, 27, 410, 151, 174, 29, 287, 11, 304, 347, 161, 285, 437, 341, 366, 5, 225, 334, 423, 297, 47, 324, 147, 371, 38, 323, 53, 184, 344, 378, 162, 394, 414, 106, 110, 426, 346, 130, 264, 432, 179, 440, 191, 94, 189, 362, 316, 220, 175, 421, 39, 188, 19, 401, 389, 419, 309, 224, 109, 95, 197, 97, 31, 244, 135, 351, 133, 352, 402, 63, 405, 23, 81, 305, 396, 330, 266, 303, 294, 43, 62, 364, 295, 168, 181, 165, 114, 89, 66, 250, 248, 278, 13, 430, 25, 416, 34, 398, 199, 84, 56, 381, 76, 35, 269, 291, 404, 228, 185, 241, 453, 211, 160, 15, 209, 173, 420, 326, 286, 64, 273, 9, 299, 382, 139, 442, 96, 372, 201, 30, 446, 386, 403, 127, 196, 69, 155, 75, 360, 231, 235, 310, 267, 255, 345, 349, 171, 126, 190, 452, 342, 57, 182, 59, 455, 439, 123, 383, 112, 59, 357, 55, 142, 214, 203, 374, 314, 170, 87, 187, 300, 355, 262, 7, 260, 268, 280, 1, 44, 298, 265, 120, 249, 91, 422, 436, 18, 108, 113, 176, 283, 245, 433, 158, 218, 122, 313, 172, 380, 365, 154, 385, 290, 388, 247, 216, 67, 98, 368, 177, 164, 289, 320, 32, 186, 400, 16, 237, 42, 431, 157, 375, 272, 406, 259, 315, 117, 77, 152, 306, 434, 72, 166, 392, 449, 230, 246, 26, 445, 232, 146, 118, 107, 61, 129, 50, 143, 427, 144, 370, 311, 336, 447, 105, 263, 256, 215, 125, 167, 240, 226, 195, 48, 194, 200, 3, 213, 354, 340, 45, 141, 102, 332, 51, 443, 33, 68].
Those papers that reference this claim without citation include [78, 212, 428, 116, 8, 223, 407, 4, 217, 261, 70, 169, 343, 153, 219, 206, 204, 2, 350, 156, 134, 22, 115, 257, 353, 279, 85, 424, 258, 367, 10, 454, 251, 275, 140, 356, 79, 52, 444, 90, 252, 111, 296, 293, 321, 358, 41, 54, 159, 369, 71, 180, 301, 74, 363, 24, 86, 253, 270, 390, 329, 333, 348, 233, 6, 124, 448, 359, 236, 243, 335, 210, 284, 21, 12, 379, 274, 202, 415, 377, 373, 271, 82, 376, 242, 183, 88, 322, 395, 239, 338, 149, 288, 408, 136, 128, 46, 163, 28, 412, 101, 328, 73, 60, 20, 103, 151, 29, 287, 11, 304, 161, 285, 341, 5, 423, 324, 371, 53, 184, 344, 162, 414, 110, 346, 264, 179, 191, 94, 189, 362, 316, 175, 188, 19, 389, 419, 109, 95, 351, 133, 23, 396, 303, 43, 364, 168, 165, 114, 250, 248, 430, 25, 416, 34, 398, 199, 56, 76, 35, 269, 291, 404, 228, 241, 453, 211, 160, 15, 420, 286, 9, 299, 139, 442, 96, 372, 30, 196, 69, 75, 360, 231, 235, 267, 171, 190, 342, 57, 439, 383, 112, 59, 357, 214, 314, 87, 187, 300, 262, 7, 260, 280, 44, 120, 249, 91, 18, 108, 113, 176, 283, 245, 122, 172, 380, 154, 385, 290, 216, 67, 164, 289, 186, 400, 16, 42, 431, 157, 375, 77, 306, 166, 449, 230, 232, 146, 118, 107, 61, 370, 311, 336, 447, 105, 263, 256, 125, 240, 3, 354, 340, 45, 141, 102, 332, 443]
Those that do so while citing only other papers that themselves never reference or argue Claim 1 include [428, 219, 204, 350, 156, 134, 22, 454, 251, 275, 356, 79, 52, 444, 90, 358, 41, 159, 369, 74, 86, 270, 333, 6, 124, 359, 243, 210, 284, 12, 379, 415, 377, 373, 395, 239, 136, 128, 328, 73, 20, 103, 29, 11, 304, 285, 5, 162, 414, 264, 191, 94, 189, 175, 419, 133, 303, 364, 165, 250, 25, 416, 398, 35, 228, 241, 453, 211, 160, 15, 286, 139, 96, 372, 30, 69, 75, 360, 231, 235, 171, 112, 59, 357, 314, 262, 7, 44, 120, 18, 113, 176, 283, 122, 172, 154, 385, 290, 216, 289, 431, 157, 375, 230, 146, 118, 107, 61, 311, 336, 3, 45, 141, 332]
All papers identified, manual screening results, and extracted quotes will be made available upon publication.